OpenStudy (anonymous):

What is the solution to the system of equations? -2x-3y+z=-6 z=6 3x-y+z=13 A.(3, 2, 6) B.(3, 2, –6) C.(3, –2, 6) D.(–3, 2, 6)

OpenStudy (johnweldon1993):

Well if we already know that z = 6 we can plug that into both the equations we have -2x - 3y + z = -6 becomes -2x - 3y + 6 = -6 subtract 6 from both sides \(\large \color \red{-2x - 3y = -12}\) And the second equation we have is 3x - y + z = 13 This will become 3x - y + 6 = 13 subtract 6 from both sides again \[\large \color \red{3x - y = 7}\] Now we just have a system of 2 equations...can you solve them?

OpenStudy (texaschic101):

you type fast johnweldon

OpenStudy (anonymous):

no, i am horrible at these. i always get the wrong answer.

OpenStudy (johnweldon1993):

My computer was lagging so it wasn't registering @texaschic101 lol And okay we'll work through it @alo03209

OpenStudy (texaschic101):

she skipped out on you

OpenStudy (johnweldon1993):

The nerve! haha

OpenStudy (texaschic101):

lol

OpenStudy (anonymous):

okay.

OpenStudy (texaschic101):

-2x - 3y = -12 3x - y = 7 --->(-3) ------------------ -2x - 3y = -12 -9x + 3y = -21 (result of multiplying by -3) ------------------add -11x = - 33 x = 3 now sub 3 in for x 3x - y = 7 3(3) - y = 7 9 - y = 7 -y = -2 y = 2 check... 3x - y = 7 3(3) - 2 = 7 9 - 2 = 7 7 = 7 (correct) x = 3, y = 2, z = 6 any questions ?

OpenStudy (anonymous):

pretty self explanitory , thanks.

OpenStudy (texaschic101):

glad to help :)